OpenStudy (anonymous):

The two lines 3x-y-1=0 and x+3y-5=0 are: a) perpendicular b) parallel c) confounded d) collinear

OpenStudy (irishboy123):

you might wish to compare their slopes re-writing each as \(y = ...\) would help you do that

OpenStudy (anonymous):

i got x=3/5. and y=4/5. but now no idea how to continue

OpenStudy (sepeario):

Is confounded points even a thing?

OpenStudy (anonymous):

the right answer is a- but how to know if 2 lines are perp/paral/conf ?

OpenStudy (irishboy123):

what you have done is found *incorrectly* the intersection point. plug your answer back into both equations to see. as i said, if you re-write both equations in \(y = mx + c\) form, you will see m, ie the slope, for each. we can go from there. [and i know that n some countries they use \(y = ax + b\) or \(y = mx + b\) but do that first....]

OpenStudy (irishboy123):

@Sepeario i don't think so. i actually looked it up :p

OpenStudy (anonymous):

oki so i got the first y1=3x-1 and y2=(-x/3)-(5/3) what is the next step?

OpenStudy (irishboy123):

y2 has an error in it but this is the missing bit of information you need when you multiply the slopes of lines that are perpendicular, you get -1 https://gyazo.com/1fa13ef12ef26796eab06a832615537a

OpenStudy (anonymous):

ohhh thank you so much you were so helpful. A little more question, how to know if they are parallel?

OpenStudy (phi):

FYI, you should get the equations \[ y = 3x -1 \\y= -\frac{1}{3} x + \frac{5}{3} \] compare the slopes: if they are the same the lines are parallel (unless you have the exact same equation for both, in which case they are the same line) if the slopes are *negative reciprocals*, the lines are perpendicular (form a right angle where they cross) negative reciprocal means "flip" and multiply by -1 for examples: 2/3 and -3/2 (first flip 2/3 to get 3/2 , then negate to get -3/2) or 2 and -1/2 : flip 2 (think of it as 2/1) to get 1/2 , then negate: -1/2 or -1/4 and 4 etc the other test is if you multiply them you get -1 as Irish explained up above